OpenStudy (anonymous):

Solve the equation using the Zero-Product Property. (x-2)(x+7)=0

OpenStudy (anonymous):

use the distribute property

OpenStudy (phi):

look at (x-2)(x+7)=0 if x were 2, what would you get ? replace x with 2 to find out: (2-2) ( 2+7) = ??

OpenStudy (anonymous):

x-2=0 , x+7=0 x=0+2 , x=0+7 x=2 , x=7

OpenStudy (anonymous):

thanks i get it now :)

OpenStudy (phi):

if this is not clear, you should study it, because you have to know this idea to solve quadratics